jagr2713 (jagr2713):

Trick Question : Robin takes the underground train to work and uses an escalator at the railway station. If Robin runs up 6 steps of the escalator, then it takes him 21.0 seconds to reach the top of the escalator. If he runs up 13 steps of the escalator, then it takes him only 10.5 seconds to reach the top. Explain :D

jagr2713 (jagr2713):

@sleepyjess @Nnesha

jagr2713 (jagr2713):

@campbell_st

jagr2713 (jagr2713):

@Compassionate @EclipsedStar

OpenStudy (jdoe0001):

well... there is really no question is mostly an statement about Robin working himself up, up and down the escalator it seems I'd suggest to Robin to either get the elevator,or just to relax for a while

pooja195 (pooja195):

Total steps = 13 + 21.0 / 1.5 = 27.

Nnesha (nnesha):

LOLOL

pooja195 (pooja195):

LOL xD

jagr2713 (jagr2713):

ok @jdoe0001 and @pooja195 Nope O_O

pooja195 (pooja195):

really?

jagr2713 (jagr2713):

Yep D: you were fine yesterday D:

pooja195 (pooja195):

7 + 30.0 / 1.5 = 27 27+27=40

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @jagr2713 ok @jdoe0001 and @pooja195 Nope O_O \(\color{blue}{\text{End of Quote}}\) lol so you are saying he can't take elevator :P

jagr2713 (jagr2713):

Nope D;

OpenStudy (sleepyjess):

27 + 27 = 40?????

jagr2713 (jagr2713):

Nope D:

pooja195 (pooja195):

i dunno this one >_< jess my two equations combined

OpenStudy (sleepyjess):

How does 27 + 27 equal 40 though?

jagr2713 (jagr2713):

LOLOLOL

pooja195 (pooja195):

lol 54 XD

OpenStudy (sleepyjess):

lol

jagr2713 (jagr2713):

you guys give up ?

pooja195 (pooja195):

i do :P

jigglypuff314 (jigglypuff314):

rate = difference of seconds divided by difference of steps = 21-10.5 / 13-6 = 10.5 / 7 = 1.5 1.5 seconds per step 21.0 seconds divided by 1.5 = 14 steps + 6 starting steps = total of 20 steps 20 steps at 1.5 rate = 30 seconds should've commented as sleepy, I'm mubbling

OpenStudy (sleepyjess):

Emily was typing...

jagr2713 (jagr2713):

YES @jigglypuff314

OpenStudy (sleepyjess):

mubbline? o.O

jigglypuff314 (jigglypuff314):

your question wasn't very clear http://www.braingle.com/brainteasers/teaser.php?op=2&id=5079&comm=0

OpenStudy (sleepyjess):

*mubbling

pooja195 (pooja195):

LOL

jigglypuff314 (jigglypuff314):

mummbling

OpenStudy (sleepyjess):

haha

jagr2713 (jagr2713):

thats a different question @jigglypuff314

jigglypuff314 (jigglypuff314):

well idk what you were asking :P so I tried answering all of them xP

OpenStudy (sleepyjess):

xDDD

jagr2713 (jagr2713):

yep :D i wanted to change the numbers so you dont find it on google.

jigglypuff314 (jigglypuff314):

but I did ;) #googleisgod

jagr2713 (jagr2713):

lol